State the degrees of freedom error in each of the following tests.

(a) A consultant measures job satisfaction in a sample of 16 supervisors, 16 managers, and 16 executives at a local firm.


(b) A researcher tests how nervous public speakers get in front of a small, medium, or large audience. Eleven participants are randomly assigned to each group.


(c) A high school counselor has 12 students in each of four classes rate how much they like their teacher.
Answer #1

Solution :

a)

n = 16*3 = 48, k = 3

Hence,

Error df = 48 - 3 = 45

b)

n = 11*3 = 33, k = 3

Hence,

Error df = 33 - 3 = 30

c)

n = 4*12 = 48, k = 4

Hence,

Error df = 48 - 4 = 44


Let me know in the comment section if anything is not clear. I will reply ASAP!

If you like the answer, please give a thumbs-up. This will be quite encouraging for me.Thank-you!